OpenStudy (anonymous):

The distance in feet that a free-falling body falls in each second, starting with the first second, is given by the sequence 21, 63, 105, 147, . . . . Find the distance that the body falls in the 7th second.

OpenStudy (anonymous):

it is an arithmatic progression a= 21 d= 42 so a5=a+(4)*d =21 + 168 = 189